OpenStudy (anonymous):

Tanya has a balance of $1700 on a credit card with an APR of 24.2%, compounded monthly. About how much will she save in interest over the course of a year if she transfers her balance to a credit card with an APR of 10.8%, compounded monthly? (Assume that Tanya will make no payments or new purchases during the year, and ignore any possible late payment fees.) A. $202.61 B. $152.51 C. $314.28 D. $267.28 @lulu22 @ganeshie8 @helpme1.2 @live10000000 @John_ES @rrp240

OpenStudy (john_es):

\[1700 \left(1+\frac{24.2}{1200}\right)^{12}-1700 \left(1+\frac{10.8}{1200}\right)^{12}=267.28\]

OpenStudy (anonymous):

APEX: $267.28 ~~~ EDDIE DARK CONFIRMED YO! ;)